Circular No. 7596

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ac IN NGC 3504 S. Beckmann and W. D. Li, University of California at Berkeley, on be-

half of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova (mag ∼ 18.2) on unfiltered CCD images taken on Mar. 12.4 and 13.3 UT with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2001ac is located at α = 11^{h}03^{m}15^{s}.37, δ = +27^{o}58′29′′.5 (equinox 2000.0), which is 55′′.3 east and 9′′.1 north of the nucleus of NGC 3504. A KAIT image taken on Feb. 4.4 UT showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001ad IN NGC 6373 Li also forwards the report by X. Y. Dong, Y. L. Qiu, and J. Y. Hu, on

behalf of the Beijing Astronomical Observatory (BAO) Supernova Survey, of the discovery of an apparent supernova (mag ∼ 17.4) on an unfiltered CCD image taken with the BAO 0.6-m telescope on Mar. 11.8 UT. SN 2001ad is located at α = 17^{h}24^{m}02^{s}.45, δ = +58^{o}59′52′′.2 (equinox 2000.0), which is 43′′.2 west and 10′′.3 north of the nucleus of NGC 6373. SN 2001ad was confirmed at mag 18.4 on an unfiltered KAIT image taken on Mar. 13.5, from which Li measured the position end figures to be 02^{s}.40, 52′′.0, which is 44′′.3 west and 9′′.6 north of the galaxy nucleus. Images taken with the BAO 0.6-m telescope on Mar. 8.8 (limiting mag ∼ 17.5) and with KAIT on 2000 Sept. 9 (limiting mag ∼ 19.5) showed nothing at this position.

COMET C/1999 T1 (McNAUGHT-HARTLEY) E. A. Bergin, Harvard-Smithsonian Center for Astrophysics (CfA); D.

A. Neufeld, Johns Hopkins University; and S. C. Kleiner, Z. Wang, and G. J. Melnick, CfA, write: “The 1(10)–1(01) transition of water vapor near 557 GHz was detected toward comet C/1999 T1 by the Submillimeter Wave Astronomy Satellite. During the periods Feb. 2.01–11.06 and 23.01–28.95 UT, the average integrated antenna temperatures were 0.58 ± 0.02 and 0.39 ± 0.03 K km s^{−1}, respectively, within a 3′.3 × 4′.5 (FWHM) elliptical beam. For a spherical outflow model with an assumed water lifetime of 7.3 × 10^{4} s and an assumed water ortho-para ratio of 3, the inferred total water production rates (× 10^{28} molecules/s) are 5.7 and 4.4, respectively.”

2001 March 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green